# Extrinsic contribution to bosonic thermal Hall transport

### v1

@@ II.1 Reminders about the clean case
We start from the general formulation introduced in Ref. Mangeolle et al. , of a free boson system described by $2N$ hermitian fields $\Phi_{a}(r)$, $r\in\mathbb{R}^{d}$, $a=1,..,2N$, with commutation relation $[{\Phi}_{a}(r_{1}),{\Phi}_{b}(r_{2})]\equiv\hbar\,\hat{\sf\Gamma}_{ab}(r_{1},r_{2})$ a $2N\times 2N$ matrix with pure imaginary $c$-number entries. The free theory without disorder is defined by an arbitrary quadratic hamiltonian, $\textrm{H}=\frac{1}{2}{\rm Tr}\left[\Phi\circ\hat{\sf H}\circ\Phi\right]$ where the convolution $\circ$ and the trace Tr run over both flavor and spatial indices, explicitly
@@ II.2 Introducing disorder
$\partial_{t}\hat{\sf F}$ $=-\frac{i}{\hbar}\big{(}(\hat{\sf K}+\hat{\sf V})\circ\hat{\sf F}-\hat{\sf F}\circ(\hat{\sf K}+\hat{\sf V})^{\dagger}\big{)},$ (9)
@@
The new ingredient here is $\hat{V}$, accounting for disorder. Generally, $\hat{V}_{ab}(r_{1},r_{2})$ has an arbitrary matrix structure, only constrained by the same symmetry properties as $\hat{\sf H}_{ab}(r_{1},r_{2})$. In the existing treatments of disorder in (fermionic) anomalous Hall effects Adams and Blount ; Fivaz ; Nozières and Lewiner ; Sinitsyn et al. (2005, 2006, 2007); König and Levchenko , a diagonal scalar potential $\hat{V}_{ab}(r_{1},r_{2})=\delta_{a,b}\,\delta(r_{1}-r_{2})\,V(r)$ is used, which corresponds effectively to an on-site shift of the electronic chemical potential  [^3] In König and Levchenko an extra valley index is introduced, but the $(a,b)$ and $(r_{1},r_{2})$ indices remain decoupled, the spatial dependence is still of the $\delta(r_{1}-r_{2})$ type, and within one valley disorder is still $\hat{V}\propto\text{\textbb{1}}$.; in the following we will refer to this as the “scalar fermionic case”. For real bosons this would be unreasonable in general: because of the relation $\hat{\sf V}=\hbar\,\hat{\sf\Gamma}\circ\hat{V}$, an on-site chemical potential (i.e. a term $\hat{\sf V}\propto\text{\textbb{1}}$) corresponds in general to a nontrivial and fine-tuned matrix structure for $\hat{V}$. We also drop the assumption $\hat{V}_{ab}(r_{1},r_{2})\propto\delta(r_{1}-r_{2})$ and instead consider a more general spatial dependence. This allows us to describe structural disorder, bond disorder, impurities interacting or mediating interactions at longer range, and more generally any local disturbance of momentum-dependent features of the clean theory (as opposed to just an on-site scalar potential).
@@
Explicitly, we consider $\hat{V}_{ab}(r_{1},r_{2})$ a gaussian random variable with zero average, $\langle\hat{V}_{ab}(r_{1},r_{2})\rangle=0$, interpreted as a deterministic function of $(r_{1}-r_{2})$ centered at a position $\tfrac{1}{2}(r_{1}+r_{2})$ and random with respect to the latter. This means, mathematically,
$\langle\hat{V}_{ab}(\lambda_{1},\lambda_{2})\hat{V}_{cd}(\lambda_{3},\lambda_{4})\rangle$
$=\hat{W}_{ab}(\lambda_{1}-\lambda_{2})\hat{W}_{cd}(\lambda_{3}-\lambda_{4})\;w(\tfrac{\lambda_{1}+\lambda_{2}}{2}-\tfrac{\lambda_{3}+\lambda_{4}}{2}),$ (10)
where the first two factors describe the (deterministic) matrix structure of single impurities, and the function $w$ describes the impurities’ spatial correlation – we assume $w^{*}(r)=w(r)=w(-r)$. We choose the normalization $w(0)=n_{\rm int}$, the impurity density, to remove ambiguity. For future convenience, we also define $\hat{\sf W}=\hat{\mathsf{\Gamma}}\circ\hat{W}$. We note that the general disorder we consider here is similar to that of Culcer et al. , where a homogeneous problem is considered in a full momentum representation – at odds with our approach of an inhomogeneous problem in a phase-space representation, allowing for a semiclassical identification of wavepacket coordinate shifts.
@@ II.3 Equal-time formalism
with the shorthand $[{\sf K}_{\rm d}\,\overset{\circ}{,}\,{\sf F}_{\rm d}]={\sf K}_{\rm d}\circ{\sf F}_{\rm d}-{\sf F}_{\rm d}\circ{\sf K}_{\rm d}$. Note that $\mathsf{V}_{\rm d}$ is a random variable, and consequently ${\sf F}_{\rm d}$, as a solution of Eq.(11), is also random: to extract physical quantities, a disorder average of Eq.(11) must be taken.
@@
×$\qquad\times\left[\mathsf{S}^{-1}(z_{1},\lambda_{1})\mathsf{W}(\lambda_{1},\lambda_{2})\mathsf{S}(\lambda_{2},z_{2})\right]^{s}_{n,n^{\prime}}\;\left[\mathsf{S}^{-1}(z_{3},\lambda_{3})\mathsf{W}(\lambda_{3},\lambda_{4})\mathsf{S}(\lambda_{4},z_{4})\right]^{s^{\prime}}_{n^{\prime},n},$
where the integral is over all four $\lambda_{i}$, $i=1,\dots,4$ and the upper index $s,s^{\prime}=-$ (resp. $+$) on an operator means hermitian conjugation (resp. absence thereof).
[…]
@@ III.1 Phase space formulation and semiclassical approximations
First, our semiclassical derivation relies on an expansion in powers of $\hbar$. Scattering of bosons on disorder affects the phase-space coordinates $X$ and $p$ of the wavepackets, which by convention are $O(\hbar^{0})$ in the free case. Upon scattering, the particle momentum is modified, $p\rightarrow p^{\prime}$, by an amount $p^{\prime}-p=O(\hbar^{0})$ governed by the Fourier-transformed disorder correlation $\tilde{w}(k)=\int_{\xi}e^{-ik\xi/\hbar}w(\xi)$: since the latter can hardly be known with sufficient accuracy, any $O(\hbar^{1})$ semiclassical correction is irrelevant, and we neglect scattering shift effects on the momentum of excitations. Meanwhile, we focus on the position shift effects induced by disorder which will occur at order $O(\hbar^{1})$ – in other words, we consider semiclassical corrections on the $X$ but not the $p$ variable of ${\sf F}_{\rm d}(X,p),{\sf K}_{\rm d}(X,p)$ in our derivation.
@@
where $\mathcal{P}$ stands for principal value. Physically, this amounts to assuming destructive intereferences between all contributions in Eq.(15) at times $\tau$ earlier than $t$, yielding an equal-time differential equation for the time evolution of the density matrix. Equivalently, the $\hbar\rightarrow 0$ limit selects the classical path in the functional integral representation of time evolution. In the following we also discard the principal value term, which accounts for the perturbative effect of disorder on the bosonic quasiparticle energies, and keep only the dirac delta term, which accounts for elastic scattering of bosonic waves.
@@
Contrary to the previous two approximations, Eq.(III.1) is in principle an uncontrolled assumption, which amounts to neglecting variations of the moduli of eigenvectors, focussing on their phase variations. { In the scalar fermionic case, because $\mathsf{W}\propto\text{\textbb{1}}$ and $\mathsf{S}^{\dagger}\circ\mathsf{S}=\text{\textbb{1}}$, Eq.(III.1) is strictly equivalent to $k\ll p{}^{(}{}^{\prime}{}^{)}$ ; in our general bosonic case there is no such equivalence. } In the wavepacket interpretation of phase-space quantum dynamics, this means we are focussing on the position and momentum of particle wavepackets (contained in the phase), neglecting variations of their extension or spreading (contained in the amplitude): this is again a semiclassical approximation in the traditional sense Gutzwiller (1990). Note also that in Eqs.(III.1),(19) we dropped the $X$ variable of all quantities ($\mathsf{S},\mathsf{M},\delta\mathsf{r}$, etc): while these may indeed vary spatially in our inhomogeneous theory, this dependence is irrelevant to the order of $\hbar$ considered here, and we will keep omitting the $X$ variable in the following wherever this cannot possibly cause any confusion.
@@
×$\qquad\times\bigg{\{}(\mathsf{F}_{\rm d})_{n,n}(X,p,t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n^{\prime},n}(p^{\prime},p)\right]-(\mathsf{F}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta\mathsf{r}_{n,n^{\prime}}^{-}(p,p^{\prime}),p^{\prime},t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})^{*}\right]\bigg{\}},$
where $[\mathsf{O}_{1}\overset{\star}{,}\mathsf{O}_{2}]=\mathsf{O}_{1}\star\mathsf{O}_{2}-\mathsf{O}_{2}\star\mathsf{O}_{1}$ is the Moyal bracket, and $\delta\mathsf{r}_{n,n^{\prime}}^{-}(p,p^{\prime})=\tfrac{1}{2}[\delta\mathsf{r}_{n,n^{\prime}}(p,p^{\prime})-\delta\mathsf{r}_{n^{\prime},n}(p^{\prime},p)]$ appears as a coordinate shift. We note that $\tilde{w}^{*}(k)=\tilde{w}(k)=\tilde{w}(-k)$.
@@ III.2 Gauge-invariant coordinates and shift
$-ϑ\displaystyle\mathsf{S}\rightarrow\mathsf{S}\star e^{i\stheta}$ (21)
at any point $(X,p)$ in phase space ($-ϑ\stheta$ a diagonal matrix of phases, function of $X$ and $p$). As explained e.g. in Mangeolle et al. , ${\sf K}_{\rm d}(X,p)$ and ${\sf F}_{\rm d}(X,p)$ themselves are not invariant under such a gauge transformation Eq.(21). Instead, $\underline{\sf K}_{\rm d}(X,p)={\sf K}_{\rm d}(\underline{X},\underline{p})$ and $\underline{\sf F}_{\rm d}(X,p)={\sf F}_{\rm d}(\underline{X},\underline{p})$, functions of the “kinetic” coordinates
@@
The above results hold for a general inhomogeneous theory. In the following, we focus on physical problems where the mixed Berry curvatures $\mathsf{\Omega}_{X,X},\mathsf{\Omega}_{X,p},\mathsf{\Omega}_{p,X}$ and the real-space Berry connection $\mathsf{A}_{X_{\mu}}$ play no role (which includes not only homogeneous theories but also, for instance, the “gravitational potential” or “separable” case Luttinger (1964); Mangeolle et al. ).
@@
×$\times\bigg{\{}(\underline{\mathsf{F}}_{\rm d})_{n,n}(X,p,t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n^{\prime},n}(p^{\prime},p)\right]-(\underline{\mathsf{F}}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta{\sf X}_{n,n^{\prime}}(p,p^{\prime}),p^{\prime},t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})^{*}\right]\bigg{\}}.$
@@ III.3 Corrected kinetic equation: practical implementation and further assumptions
First we will replace $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$ by its modulus: while it is clear that this does not affect the semiclassical physics of coordinate shifts but merely modifies the scattering rates, whose precise form ultimately does not affect our final results, it is also possible to justify this assumption more rigorously based on a small-momentum expansion (see Appendix C.3). Second, contrary to the scalar fermionic case, where it is always true,
@@
×$\times\big{|}\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\big{|}^{2}\times\delta F_{n,n^{\prime}}(p,p^{\prime}),$ (29)
@@
and the energy conservation delta-function is the same as in Eq.(27). Besides, in the following we neglect $n\neq n^{\prime}$ contibutions to the collision integral. An argument based on matrix elements $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$ is given in App.C.2, but here we simply notice that for non-degenerate bands and small deviations $p^{\prime}-p$, the energy conservation delta strongly favors $n^{\prime}=n$ terms, which we henceforth focus on.

### v2

@@ II.1 Reminders about the clean case
We start from the general formulation introduced in Ref. Mangeolle et al. , of a free boson system described by $2N$ hermitian fields $\Phi_{a}(r)$, $r\in\mathbb{R}^{d}$, $a=1,..,2N$, with commutation relation ${\Phi}_{a}(r_{1}),{\Phi}_{b}(r_{2})]\equiv\hbar\,\hat{\sf\Gamma}_{ab}(r_{1},r_{2})$ a $2N\times 2N$ matrix with pure imaginary $c$-number entries. The free theory without disorder is defined by an arbitrary quadratic hamiltonian, $\textrm{H}=\frac{1}{2}{\rm Tr}\left[\Phi\circ\hat{\sf H}\circ\Phi\right]$ where the convolution $\circ$ and the trace Tr run over both flavor and spatial indices, explicitly
@@ II.2 Introducing disorder
$\partial_{t}\hat{\sf F}$ $),\displaystyle=-\frac{i}{\hbar}\big((\hat{\sf K}+\hat{\sf V})\circ\hat{\sf F}-\hat{\sf F}\circ(\hat{\sf K}+\hat{\sf V}$$^{\dagger}\big)$, (9)
@@
The new ingredient here is $\hat{V}$, accounting for disorder. Generally, $\hat{V}_{ab}(r_{1},r_{2})$ has an arbitrary matrix structure, only constrained by the same symmetry properties as $\hat{\sf H}_{ab}(r_{1},r_{2})$. In the existing treatments of disorder in (fermionic) anomalous Hall effects Adams and Blount ; Fivaz ; Nozières and Lewiner ; Sinitsyn et al. (2005, 2006, 2007); König and Levchenko , a diagonal scalar potential $\hat{V}_{ab}(r_{1},r_{2})=\delta_{a,b}\,\delta(r_{1}-r_{2})\,V(r)$ is used, which corresponds effectively to an on-site shift of the electronic chemical potential  [^3] In König and Levchenko an extra valley index is introduced, but the $(a,b)$ and $(r_{1},r_{2})$ indices remain decoupled, the spatial dependence is still of the $\delta(r_{1}-r_{2})$ type, and within one valley disorder is still $\hat{V}\propto\text{\textbb{1}}$.; in the following we will refer to this as the “case of fermions in a random scalar potential”. For real bosons this would be unreasonable in general: because of the relation $\hat{\sf V}=\hbar\,\hat{\sf\Gamma}\circ\hat{V}$, an on-site chemical potential (i.e. a term $\hat{\sf V}\propto\text{\textbb{1}}$) corresponds in general to a nontrivial and fine-tuned matrix structure for $\hat{V}$. We also drop the assumption $\hat{V}_{ab}(r_{1},r_{2})\propto\delta(r_{1}-r_{2})$ and instead consider a more general spatial dependence. This allows us to describe structural disorder, bond disorder, impurities interacting or mediating interactions at longer range, and more generally any local disturbance of momentum-dependent features of the clean theory (as opposed to just an on-site scalar potential).
Explicitly, we consider $\hat{V}_{ab}(r_{1},r_{2})$ a gaussian random variable with zero average, $\langle\hat{V}_{ab}(r_{1},r_{2})\rangle=0$, interpreted as a fixed (i.e. non-random) function of $(r_{1}-r_{2})$ centered at a position $\tfrac{1}{2}(r_{1}+r_{2})$ which is random. This means, mathematically,
$\langle\hat{V}_{ab}(r_{1},r_{2})\hat{V}_{cd}(r_{3},r_{4})\rangle$
$=\hat{W}_{ab}(r_{1}-r_{2})\hat{W}_{cd}(r_{3}-r_{4})\;w(\tfrac{r_{1}+r_{2}}{2}-\tfrac{r_{3}+r_{4}}{2}),$ (10)
where the first two factors describe the fixed matrix structure of single impurities, and the function $w$ describes the impurities’ spatial correlation – we assume $w^{*}(r-r^{\prime})=w(r-r^{\prime})=w(r^{\prime}-r)$. We choose the normalization $w(0)=n_{\rm imp}$, the impurity density, to remove ambiguity. For future convenience, we also define $\hat{\sf W}=\hat{\mathsf{\Gamma}}\circ\hat{W}$. We note that the general disorder we consider here is similar to that of Culcer et al. , where a homogeneous problem is considered in a full momentum representation – at odds with our approach of an inhomogeneous problem in a phase-space representation, allowing for a semiclassical identification of wavepacket coordinate shifts.
We provide more intuition on Eq.(II.2), together with a concrete recipe to compute $\hat{\sf W}$, in the more convenient language of phase space coordinates in Sec.III.5. For now, it suffices to say that $\hat{W}_{ab}(x-x^{\prime})$, the “disorder matrix”, accounts for the local matrix structure of an impurity in the basis of the field components $\Phi_{a}(x),\Phi_{b}(x^{\prime})$.
@@ II.3 Equal-time formalism
with the shorthand ${\sf K}_{\rm d}\,\overset{\circ}{,}\,{\sf F}_{\rm d}]={\sf K}_{\rm d}\circ{\sf F}_{\rm d}-{\sf F}_{\rm d}\circ{\sf K}_{\rm d}$. Note that $\mathsf{V}_{\rm d}$ is a random variable, and consequently ${\sf F}_{\rm d}$, as a solution of Eq.(11), is also random: to extract physical quantities, a disorder average of Eq.(11) must be taken.
@@
$\qquad\times\left[\mathsf{S}^{-1}(z_{1},\lambda_{1})\mathsf{W}(\lambda_{1},\lambda_{2})\mathsf{S}(\lambda_{2},z_{2})\right]^{s}_{n,n^{\prime}}\;\left[\mathsf{S}^{-1}(z_{3},\lambda_{3})\mathsf{W}(\lambda_{3},\lambda_{4})\mathsf{S}(\lambda_{4},z_{4})\right]^{s^{\prime}}_{n^{\prime},n},$
where the integral is over all four $\lambda_{i}$, $i=1,\dots,4$ and the upper index $s,s^{\prime}=$- (resp. ++) on an operator means hermitian conjugation (resp. absence thereof).
Taking $t^{\prime}\rightarrow$-$\infty$, Eq.(15) gives ${\sf F}_{\rm d}(t)$ as a function of ${\sf F}_{\rm d}(\tau)$ at all earlier times $\tau<t$. Considering ${\sf F}_{\rm d}(-\infty)$ to be disorder-independent, then ${\sf F}_{\rm d}(\tau)$ on the right-hand side of Eq.(15), as a solution to the equation at an earlier time, is no longer a random variable but a deterministic function. Therefore we can drop the disorder average $\langle\cdot\rangle$ on the left-hand side, and henceforth treat Eq.(15) as a deterministic integro-differential equation for ${\sf F}_{\rm d}(t)$.
@@ III.1 Phase space formulation and semiclassical approximations
First, our semiclassical derivation relies on an expansion in powers of $\hbar$. Throughout the paper, this is referred to as the semiclassical expansion foo . Scattering of bosons on disorder affects the phase-space coordinates $X$ and $p$ of the wavepackets, which by convention are $O(\hbar^{0})$ in the free case. Upon scattering, the particle momentum is modified, $p\rightarrow p^{\prime}$, by an amount $p^{\prime}-p=O(\hbar^{0})$ governed by the Fourier-transformed disorder correlation $\tilde{w}(k)=\int_{\xi}e^{-ik\xi/\hbar}w(\xi)$: since the latter can hardly be known with sufficient accuracy, any $O(\hbar^{1})$ semiclassical correction is irrelevant, and we neglect scattering shift effects on the momentum of excitations. Meanwhile, we focus on the position shift effects induced by disorder which will occur at order $O(\hbar^{1})$ – in other words, we consider semiclassical corrections on the $X$ but not the $p$ variable of ${\sf F}_{\rm d}(X,p),{\sf K}_{\rm d}(X,p)$ in our derivation.
@@
where $\mathcal{P}$ stands for principal value. Physically, this amounts to assuming destructive intereferences between all contributions in Eq.(15) at times $\tau$ earlier than $t$, yielding an equal-time differential equation for the time evolution of the density matrix. Equivalently, the $\hbar\rightarrow 0$ limit selects the classical path in the functional integral representation of time evolution. In the following we also discard the principal value term, which accounts for the perturbative foo effect of disorder on the bosonic quasiparticle energies, and keep only the dirac delta term, which accounts for elastic scattering of bosonic waves.
@@
Contrary to the previous two approximations, Eq.(III.1) is in principle an uncontrolled assumption, which amounts to neglecting variations of the moduli of eigenvectors, focussing on their phase variations. In the wavepacket interpretation of phase-space quantum dynamics, this means we are focussing on the position and momentum of particle wavepackets (contained in the phase), neglecting variations of their extension or spreading (contained in the amplitude): this is again a semiclassical approximation in the traditional sense Gutzwiller (1990).
{ In the case of fermions in a random scalar potential, because $\mathsf{W}\propto\text{\textbb{1}}$ and $\mathsf{S}^{\dagger}\circ\mathsf{S}=\text{\textbb{1}}$, Eq.(III.1) is strictly equivalent to $k\ll p{}^{(}{}^{\prime}{}^{)}$ ; in our general bosonic case there is no such equivalence. }
Note also that in Eqs.(III.1),(19) we dropped the $X$ variable of all quantities ($\mathsf{S},\mathsf{M},\delta\mathsf{r}$, etc): while these may indeed vary spatially in our inhomogeneous theory, this dependence is irrelevant to the order of $\hbar$ considered here, and we will keep omitting the $X$ variable in the following wherever this cannot possibly cause any confusion.
@@
$\qquad\times\bigg\{(\mathsf{F}_{\rm d})_{n,n}(X,p,t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n^{\prime},n}(p^{\prime},p)\right]-(\mathsf{F}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta\mathsf{r}_{n,n^{\prime}}^{-}(p,p^{\prime}),p^{\prime},t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})^{*}\right]\bigg\},$
where $\mathsf{O}_{1}\overset{\star}{,}\mathsf{O}_{2}]=\mathsf{O}_{1}\star\mathsf{O}_{2}-\mathsf{O}_{2}\star\mathsf{O}_{1}$ is the Moyal bracket, and $\delta\mathsf{r}_{n,n^{\prime}}^{-}(p,p^{\prime})=\tfrac{1}{2}[\delta\mathsf{r}_{n,n^{\prime}}(p,p^{\prime})-\delta\mathsf{r}_{n^{\prime},n}(p^{\prime},p)]$ appears as a coordinate shift. We note that $\tilde{w}^{*}(k)=\tilde{w}(k)=\tilde{w}(-k)$.
@@ III.2 Gauge-invariant coordinates and shift
[…]
@@
The above results hold for a general inhomogeneous theory. In the following, we focus on physical problems where the “magnetic field” $\mathsf{\Omega}_{X,X}$, the mixed Berry curvatures $\mathsf{\Omega}_{X,p},\mathsf{\Omega}_{p,X}$, and the real-space Berry connection $\mathsf{A}_{X_{\mu}}$ play no role (which includes not only homogeneous theories but also, for instance, the “gravitational potential” or “separable” case Luttinger (1964); Mangeolle et al. ).
@@
$\times\bigg\{(\underline{\mathsf{F}}_{\rm d})_{n,n}(X,p,t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n^{\prime},n}(p^{\prime},p)\right]-(\underline{\mathsf{F}}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta{\sf X}_{n,n^{\prime}}(p,p^{\prime}),p^{\prime},t)\;{\rm Re}\left[\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})^{*}\right]\bigg\}.$
@@ III.3 Corrected kinetic equation: practical implementation and further assumptions
First we will replace $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$ by its modulus: while it is clear that this does not affect the semiclassical physics of coordinate shifts but merely modifies the scattering rates, whose precise form ultimately does not affect our final results, it is also possible to justify this assumption more rigorously based on a small-momentum expansion (see Appendix C.3). Second, contrary to the case of fermions in a random scalar potential, where it is always true,
@@
$\times\big|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\big|^{2}\times\delta F_{n,n^{\prime}}(p,p^{\prime}),$ (29)
@@
and the energy conservation delta-function is the same as in Eq.(27). Besides, in the following we neglect $n\neq n^{\prime}$ contributions to the collision integral. An argument based on the relative magnitude of matrix elements $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$ is given in App.C.2, but here we simply notice that for non-degenerate bands and small deviations $p^{\prime}-p$, the energy conservation delta strongly favors $n^{\prime}=n$ terms, which we henceforth focus on.